OpenStudy (anonymous):

Can someone please help i'm stuck on one part!! h(x)= 5sqrtf(x)^4

OpenStudy (tkhunny):

What is to be done?

OpenStudy (anonymous):

ok I have this so \[1/5(f(x)^4)^{1/5}(f(x)^4)'\]

OpenStudy (tkhunny):

?? We have \(g(x) = \sqrt[5]{\left[f(x)\right]^{4}}\) and we seek \(g'(x)\)?

OpenStudy (anonymous):

we have h(x) =\[\sqrt[5]{f(x)^4}\]

OpenStudy (anonymous):

find h'

OpenStudy (tkhunny):

Oh, sorry. Not sure why I switched to g(x). Ummm... What, in the name of reason, are you doing? \(h(x) = \sqrt[5]{\left[f(x)\right]^{4}} = [f(x)]^{4/5}\) \(h'(x) = \dfrac{4}{5}\left[f(x)\right]^{-1/5}\) Try to make your life simpler. No need to do anything the hardest way possible.

OpenStudy (anonymous):

ok thankyou so much because I was confused :)